A farmer measures the distance around a rect-
angular field. The length of each long side of
the rectangle is found to be 44.78 m, and the
length of each short side is found to be 22.5
m.
What is the total distance around the field?
1. Two significant figures (130 m)
2. Five significant figures (134.56 m)
3. Whole number (135 m)
4. Tenths (134.6 m)
5. Three significant figures (135 m)
6. Hundredths (134.56 m)
7. Four significant figures (134.6 m)
8. None of these

Answer by susanaramya(12)

the total distance around the field=2(44.78+22.5)m=2(67.28)m=134.56m